OpenStudy (anonymous):

A bullet is fired horizontally from the top of a building with a muzzle velocity of 150 m/s.A similar bullet dropped from the top of the same building, takes 4 sec to reach the ground. How far forward does the first bullet go before it hits the ground?Which bullet hits the ground first? What is the average speed of the dropped bullett?

OpenStudy (anonymous):

help me pleasee:(

OpenStudy (anonymous):

somebody help me lol

OpenStudy (anonymous):

1s=9,81ms/^2+....2s=2x...+. 3s=3x.....+ 4s=4x.... =....There you have the that answer. Unless the second bullet hits a mountain it won't hit the ground first:).

OpenStudy (anonymous):

im confused is the first answer 600 meters

OpenStudy (anonymous):

uuhm well first you have to figure out how far from the top until the gorund.

OpenStudy (anonymous):

answer lies in my earlier post

OpenStudy (anonymous):

i think i wrote the question wrong hold on

OpenStudy (anonymous):

A bullet is fired horizontally from the top of a building with a muzzle velocity of 500 m/s. A similar bullet dropped from the top of the same building, takes 4 sec to reach the ground. How far forward does the first bullet go before it hits the ground?Which bullet hits the ground first? What is the average speed of the dropped bullett?

OpenStudy (anonymous):

thats the right question^^^^

OpenStudy (anonymous):

well it's174,7161m around 174m for the similar bullet to hit the ground

OpenStudy (anonymous):

can u tell me the answers to all the questions lol

OpenStudy (anonymous):

just finished my own physic homework hmm :X

OpenStudy (anonymous):

Hmm...do you only seek answers or a tutorial on completing the question?

OpenStudy (anonymous):

im not here to get feathered at lol im sorry but yes i do only seek answers

OpenStudy (anonymous):

soooooo does anybody have the answers?

OpenStudy (anonymous):

\[\bullet1 v0=400m/s^2-9,81m/s^2 \bullet 2 a=9,81 m/s^2 t=4s s=v0t+(0,5)at^2\]

OpenStudy (anonymous):

well just some help got to go sleep, I am sure you can figure the rest out yourself :)

OpenStudy (anonymous):

i take chemistry lol -_____- f u c k

OpenStudy (anonymous):

hahaha I took maritim engineering, 3 months into physics and this crap is easy.

OpenStudy (anonymous):

When I dropped out from maritim engineering^^

OpenStudy (anonymous):

You might notice that the time take for both bullets to reach the ground is the same. Applying the motion formula, \[S_x=u_xt + \frac{ 1 }{ 2 }a_xt^2\], where a=0, u=500, t=4 for first bullet, you get? for the first bullet? How far? average speed for the bullet: from the first, \[S_y=\frac{ 1 }{ 2}(10)(4^2)=80m\] v=80/4=20

OpenStudy (anonymous):

a=-9,81m/s^2 and is the average gravity constant on earth.

OpenStudy (anonymous):

a is not= 0

OpenStudy (anonymous):

what are all the answers

OpenStudy (anonymous):

t=4 is wrong for that formula

OpenStudy (anonymous):

nvm ywan sleep sry

OpenStudy (anonymous):

yh, sorry about that, though g depends on what the question asks. t=4 because the only force acting on the bullet is gravity, which acts on the y-axis. Thus the time taken to go through x and y axis is the same.

OpenStudy (anonymous):

shadowys knows more than me anyways I thik:) maybe 4is right havent checked

OpenStudy (anonymous):

gn

OpenStudy (anonymous):

lol so what are the answers

OpenStudy (anonymous):

sigh. 2000m, same time and 20ms^-1. Is this something for a test? @duskwood100 lol I need someone to confirm my answers though.

OpenStudy (anonymous):

are you reading the revised question btw

OpenStudy (anonymous):

Yes. a dropped bullet take 4s to fall to the ground.

OpenStudy (anonymous):

so whats the answer to # 1 #2 amd #3

OpenStudy (anonymous):

2000m, same time and 20ms^-1

OpenStudy (anonymous):

shankssss

OpenStudy (anonymous):

lol you're welcome :)